Diffusion Morning Report 04/19/11

Ted Berg tberg at stanford.edu
Tue Apr 19 06:30:48 PDT 2011


Good Morning All,
     Looking good this morning. TEOS has high pressure after dep starts. 
TEL FS will be here today to install noise filter. Field service will be 
here today to check out SEM4160 after PM. Epi has no PH3. LTO had O2 
flow issues. Thermco 2 no TLC flow. No other issues at this time.


> 					
> 	Date: 	4/19/11 	THICKNESS-Change at 		Priority
> Tube Name/Number 	Status 				
> Tylan 1 	*up* 	** 	** 	** 	**
> Tylan 2 	*up* 	** 	** 	** 	**
> Tylan 3 	*up* 	** 	** 	** 	**
> Tylan 4 	*up* 	** 	** 	** 	**
> Tylan 5-BBr3 	*up* 	** 	** 	** 	**
> Tylan 6-POCl3 	*up* 	** 	** 	** 	**
> Tylan 7-TEOS2 	*down* 	*high pressure* 	*5.25/30* 	** 	*1*
> Tylan 8-SiGe 	*up* 	** 	*1.94/30* 	** 	**
> Tylan 9-Poly 	*up* 	** 	*.48/30* 	** 	**
> Tylan 10-Nitride 	*up* 	** 	*11.23/30* 	** 	**
> Tylan 11-LTO 	*yellow* 	*O2 flow issues* 	*16.02/30* 	** 	*2*
> Tylan 12-FGA 	*up* 	** 	** 	** 	**
> Thermco 1 	*up* 	** 	** 	** 	**
> Thermco 2 	*yellow* 	*no tlc flow* 	** 	** 	**
> Thermco 3 	*future* 	** 	** 	** 	**
> Thermco4 	*future* 	** 	** 	** 	**
> Thermco Poly1 	*up* 	*
> * 	** 	** 	**
> Thermco Poly2 	*up* 	** 	** 	** 	**
> Thermco Nitride1 	*up* 	** 	** 	** 	**
> Thermco bpsg 	*up* 	** 	** 	** 	**
> AMAT Epi 	*yellow* 	*no PH3* 	** 	** 	**
> TEL Spa 	*down* 	*installing filter* 	** 	** 	**
> Burn Box 	*up* 	** 	** 	** 	**
> Ebeam 	*down* 	*cryo compresssor dead* 	** 	** 	**
> ZYGO 	*up* 	*
> * 	** 	** 	**
> FGA2 	*up* 	** 	** 	** 	**
> SiH4 for furnaces 	** 	*A on line 446 PSI* 	*changeover working 
> well* 	** 	**
>

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20110419/6fa79304/attachment.html>


More information about the diffusion mailing list